OpenStudy (anonymous):

tickets to science museum cost $19.95 each. there is a $3 charge for each order no matter how many tickets are ordered . write an expression for the cost (in dollars) of ordering tickets. then find the total cost if you order 5 tickets.

OpenStudy (zzr0ck3r):

imagine we buy two tickets, then we pay 19.95 each so 2*19.95, but we also a pay an extra 3$ charge just for buying any tickets at all. so f(2) = 19.95(2)+3 if we buy x tickets, what will we pay?

OpenStudy (anonymous):

normal expression f(x) = 19.95(x) + 3.00 for five tickets replace x as 5 f(5)= 19.95(5)+3 which equals one hundred and three dollars

OpenStudy (anonymous):

x being the number of tickets :) hope that helped